OpenStudy (anonymous):

what is the equation of the line in standard form that passes through the point (3,4) and is parallel to the line y=-x+6

OpenStudy (anonymous):

+x on both sides y+x=6 now -y on both sides x=6-y substitue for y x=6-(4) Now graph which u can do here http://www.wolframalpha.com/input/?i=x%3d6-4&incTime=true

OpenStudy (espex):

The equation of a line in point-slope form is \((y-y_1)=m(x-x_1)\). The given equation, \(y=-x+6\), is in slope-intercept form (y=mx+b) and you can see that, -x can be rewritten as -1x, your slope (m) is -1. Any line parallel to the given will have the same slope. Your equation for a parallel line then becomes: \((y-4)=-1(x-3)\) and all you need do now is put the equation into slope-intercept form.